# We prove that if $A=\{a_{1},\dots,a_{|A|}\}\subset\{1,2,\dots,n\}$ is a Sidon set so that $|A|=n^{1/2}-L^{\prime}$, then

### v2

@@ Abstract
where $L=\max\{0,L^{\prime}\}$. As an application of this, we give easy proofs of some previously derived results. We proceed on to proving that for any $\varepsilon>0$, we have
@@
for all $n\leq N$ but at most $\mathcal{O}_{\varepsilon}\left(N^{\frac{3}{5}+\varepsilon}\right)$ exceptions.
@@
Department of Mathematics and Statistics, IISER Kolkata, Mohanpur, West Bengal 741246, India
sd19ms148@iiserkol.ac.in
key-word: Sidon set
@@ 1 Introduction
Since then, there has been an extensive amount of literature on this topic exploring a plethora of different questions about finite and infinite Sidon sets [26]. Addressing the original question of Sidon, Erdős conjectured [13] (and offered \$ 500 for a proof or disproof) that if $S(n)$ denotes the maximum possible cardinality of a Sidon subset of $[n]:=\{1,2,\dots,n\}$, then
@@
The most recent result on this was provided by Balogh et. al. [1] proving
@@
for some $\gamma\geq 0.002$. In the same spirit, Kevin O’Bryant [27] improved the constant from $0.998$ to $0.99703$, and then again to $0.98183$ with Daniel Carter and Zach Hunter [4].
Another recent improvement in the theory of Sidon sets is a solution to the Erdős Sárközy Sós problem on asymptotic Sidon bases of order $3$. After a series of developments made by Deshoulliers and Plagne [9], Kiss [22], Kiss, Rozgonyi and Sándor [23] and Cilleruelo [8], it was finally solved by Cédric Pilatte [29] using some recent results of Will Sawin [33].
However our focus is on finding a formula for the $m$-th element of a finite Sidon set. As an application of our formula, we will indicate easier proofs of the theorems recently obtained by Yuchen Ding [10, 11].
@@
$E_{I}\leq 52\;n^{1/4}\left(1+c^{1/2}\cdot n^{1/8}\right)\left(1+L^{1/2}\cdot n^{-1/8}\right)$
@@ A Sidon subset $A\subset[n]$ is called dense if
For a prime $p$, there are at least $p$ elements in $[p^{2}-1]$ such that the sums of two of these elements are different modulo $p^{2}-1$.
Theorem C (Glyn Harman, R.C. Baker, J. Pintz. [17]).
@@
Let $A\subset[n]$ be a dense Sidon set so that $|A|=n^{1/2}-L$. Then $-n^{1/4}\leq L\leq n^{21/80}$.
@@
$p_{k}\geq\sqrt{n}-n^{21/80}$
@@
Remark: Notice that if we assume $p_{k+1}-p_{k}\ll\sqrt{p_{k}}$, then an exact same line of argument will give $L\leq n^{1/4}$.
@@ 3 The $m$-th element
Consider a Sidon set $A=\{a_{1},\dots,a_{|A|}\}\subset[n]$ so that $|A|=n^{1/2}-L$ with $L\leq n^{\frac{21}{80}}$. Let $I=[a_{m}]\subset[n]$ Let $a_{m}=cn+1$ so that $c=\dfrac{a_{m}-1}{n}<1$. So, $|I|=cn$ and $\left\lvert A\cap[a_{m}]\right\rvert=m$.
By Theorem A, the number of elements in an interval $I$ of length $cn$ is
@@
and hence putting $c=\dfrac{a_{m}-1}{n}$, multiplying by $n^{1/2}$, and rearranging, we get
@@
$\sum_{m=1}^{\left\lvert A\right\rvert}a_{m}$ $=\frac{\left\lvert A\right\rvert\left(\left\lvert A\right\rvert+1\right)}{2}\sqrt{n}+\mathcal{O}\left(n^{7/8}\left\lvert A\right\rvert\right)+\mathcal{O}\left(L^{1/2}n^{3/4}\left\lvert A\right\rvert\right)$
@@
$=\frac{1}{\ell+1}\left(\sqrt{n}\right)^{\ell+1}\cdot n^{\frac{\ell}{2}}+\mathcal{O}\left(\left(\sqrt{n}\right)^{\ell}\cdot n^{\frac{4\ell+1}{8}}\right)+\mathcal{O}\left(\left(\sqrt{n}\right)^{\ell}\cdot L^{1/2}\cdot n^{\frac{2\ell+1}{4}}\right)$
@@ 4 Improvements
Theorem D (Roger Heath-Brown [20]).
@@
We finally express
@@
for all $n\leq N$ but at most $\mathcal{O}_{\varepsilon}\left(N^{\frac{3}{5}+\varepsilon}\right)$ exceptions.
@@
So, it is enough to show that $\left\lvert L\right\rvert<n^{1/4}$ happens almost always with a small number of exceptions. As noticed in the remark under Lemma 2, denoting $p_{m}$ to be the $m$-th prime, if we have
@@
then for all $n\in\left[p_{m},p_{m+1}\right]$, we have $\left\lvert L\right\rvert<n^{1/4}$.
So, the set of exceptions are precisely those $n\in\left[p_{m},p_{m+1}\right]$ with $p_{m+1}-p_{m}\geq\sqrt{p_{m}}$. The number of such $n$ is given by
$\sum_{\begin{subarray}{c}p_{n}\leq N\\ p_{n+1}-p_{n}\geq\sqrt{p_{n}}\end{subarray}}\left(p_{n+1}-p_{n}\right)\ll_{\varepsilon}N^{\frac{3}{5}+\varepsilon}$
which follows from Theorem D.
@@
for all $n\leq N$ but at most $\mathcal{O}_{\varepsilon}\left(N^{\frac{3}{5}+\varepsilon}\right)$ exceptions.
@@ 5 Acknowledgements
We would like to thank Prof. Kevin O’Bryant for bringing [27] and [4] to our attention.

### v4

@@ Abstract
where $L=\max\{0,L^{\prime}\}$. As an application of this, we give easy proofs of some previously derived results. We proceed on to proving that for a dense Sidon set $S$ and for any $\varepsilon>0$, we have
@@
for all $n\leq N$ but at most $\mathcal{O}_{\varepsilon}\left(N^{\frac{4}{5}+\varepsilon}\right)$ exceptions.
@@
Département de mathématiques et de statistique, Université de Montréal
sayan.dutta@umontreal.ca
https://sites.google.com/view/sayan-dutta-homepage
@@ 1 Introduction
Since then, there has been an extensive amount of literature on this topic exploring a plethora of different questions about finite and infinite Sidon sets [28]. Addressing the original question of Sidon, Erdős conjectured [14] (and offered \$ 500 for a proof or disproof) that if $S(n)$ denotes the maximum possible cardinality of a Sidon subset of $n]:=\{1,2,\dots,n\}$, then
@@
The most recent result on this was provided by Balogh, Füredi and Roy [2] proving
@@
for some $\gamma\geq 0.002$. In the same spirit, O’Bryant [29] improved the constant from $0.998$ to $0.99703$, and then again to $0.98183$ with Carter and Hunter [5].
Another recent improvement in the theory of Sidon sets is a solution to the Erdős Sárközy Sós problem on asymptotic Sidon bases of order $3$. After a series of developments made by Deshoulliers and Plagne [10], Kiss [23], Kiss, Rozgonyi and Sándor [24] and Cilleruelo [9], it was finally solved by Pilatte [31] using some recent results of Sawin [35].
However our focus is on finding a formula for the $m$-th element of a finite Sidon set. As an application of our formula, we will indicate easier proofs of the theorems recently obtained by Ding [11, 12].
@@
$\left\lvert E_{I}\right\rvert\leq 52\;n^{1/4}\left(1+c^{1/2}\cdot n^{1/8}\right)\left(1+L^{1/2}\cdot n^{-1/8}\right)$
@@ A Sidon subset $A\subset[n]$ is called dense if
For a prime $p$, there are at least $p$ elements in $p^{2}-1]$ such that the sums of two of these elements are different modulo $p^{2}-1$.
Theorem C (R.C. Baker, G. Harman, J. Pintz. [1]).
@@
Let $A\subset[n]$ be a dense Sidon set so that $|A|=n^{1/2}-L$. Then $n^{1/4}\leq L\ll n^{21/80}$.
@@
$p_{k}\geq\sqrt{n}-\mathcal{O}\left(n^{21/80}\right)$
so that
$\sqrt{n}-L=|A|=S_{n}\geq\sqrt{n}-\mathcal{O}\left(n^{21/80}\right)$
@@
Remark: Notice that if we assume $p_{k+1}-p_{k}\ll\sqrt{p_{k}}$, then an exact same line of argument will give $L\ll n^{1/4}$.
@@ 3 The $m$-th element
Consider a Sidon set $A=\{a_{1},\dots,a_{|A|}\}\subset[n]$ so that $|A|=n^{1/2}-L$ with $L\leq n^{\frac{21}{80}}$. Let $I=[a_{m}]\subset[n]$ . Let $a_{m}=cn$ so that $c=\dfrac{a_{m}}{n}\leq 1$. So, $|I|=cn$ and $\left\lvert A\cap[a_{m}]\right\rvert=m$.
By Theorem A, the number of elements in an interval II of length cncn is
@@
and hence putting $c=\dfrac{a_{m}}{n}$, multiplying by $n^{1/2}$, and rearranging, we get
@@
[…]
@@
$=\frac{1}{\ell+1}\left(\sqrt{n}\right)^{\ell+1}\cdot n^{\frac{\ell}{2}}+\mathcal{O}\left(\left(\sqrt{n}\right)^{\ell}\cdot n^{\frac{4\ell+3}{8}}\right)+\mathcal{O}\left(\left(\sqrt{n}\right)^{\ell}\cdot L^{1/2}\cdot n^{\frac{2\ell+1}{4}}\right)$
@@ 4 Improvements
Theorem D (R. Heath-Brown [20]).
@@
We finally express a more quantitative version of the fact that the third error term can almost always be dropped.
@@
for all $n\leq N$ but at most $\mathcal{O}_{\varepsilon}\left(N^{\frac{4}{5}+\varepsilon}\right)$ exceptions.
@@
So, it is enough to show that $\left\lvert L\right\rvert\ll n^{1/4}$ happens almost always with a small number of exceptions. As noticed in the remark under Lemma 2, denoting $p_{m}$ to be the $m$-th prime, if we have
@@
then for all $n\in\left[p_{m}^{2}-1,p_{m+1}^{2}-1\right]$, we have $\left\lvert L\right\rvert\ll n^{1/4}$.
So, the set of exceptions are among those $n\in\left[p_{m}^{2}-1,p_{m+1}^{2}-1\right]$ with $p_{m+1}-p_{m}\geq\sqrt{p_{m}}$. So, the number of exceptions is bounded by
$\sum_{\begin{subarray}{c}p_{n}\leq\sqrt{N+1}\\ p_{n+1}-p_{n}\geq\sqrt{p_{n}}\end{subarray}}\left(p_{n+1}^{2}-p_{n}^{2}\right)$ $=\sum_{\begin{subarray}{c}p_{n}\leq\sqrt{N+1}\\ p_{n+1}-p_{n}\geq\sqrt{p_{n}}\end{subarray}}\left(p_{n+1}+p_{n}\right)\left(p_{n+1}-p_{n}\right)$
$\ll\sqrt{N}\cdot\sum_{\begin{subarray}{c}p_{n}\leq\sqrt{N+1}\\ p_{n+1}-p_{n}\geq\sqrt{p_{n}}\end{subarray}}\left(p_{n+1}-p_{n}\right)$
$\ll_{\varepsilon}\sqrt{N}\cdot N^{\frac{3}{10}+\varepsilon}=N^{\frac{4}{5}+\varepsilon}$
where the first inequality uses the fact that $p_{n+1}+p_{n}\leq 3\sqrt{N+1}$ and the second inequality follows from Theorem D.
@@
for all $n\leq N$ but at most $\mathcal{O}_{\varepsilon}\left(N^{\frac{4}{5}+\varepsilon}\right)$ exceptions.
@@
It should also be noted that the recent result given in [25] improves Lemma 2 to $L\ll n^{0.26}$. Similarly, the improvement given in [22] proves that the number of exceptions in Theorem 7 is $\mathcal{O}_{\varepsilon}\left(N^{0.785+\varepsilon}\right)$.
@@ 5 Acknowledgements
We would like to thank Prof. Kevin O’Bryant, Prof. Kaisa Matomäki and Runbo Li for bringing [29], [5], [22] and [25] to our attention. We would also like to thank the anonymous referee for his/her useful comments.